Math 122  University of South Carolina
D. Meade  Spring 1998

Math 122 (Section 10) – Business Calculus

Instructor Professor Doug Meade

Office Hours: TTh 2:00– 3:30 and by prior appointment
Office:  LeConte College 300E
Phone:  777-6183

 E-mail:  meade@math.sc.edu

WWW URL  http://www.math.sc.edu/~meade/math122-S98/

Meeting Times TTh 9:30am– 10:45am, LC 412

Text Brief Calculus, preliminary edition, by Hughes-Hallet, Gleason, et al., John Wiley & Sons, 1997.

Prerequisite Qualification through placement, or a grade of C or better in Math 111 or 115.

Overview One of the main objectives of this course is for you to understand the basic concepts of calculus well enough to know when, how, and why to apply them in real-world situations and to be able to interpret and communicate the results.

To achieve this goal will require practice at a variety of numerical, graphical, and analytic methods. In particular, being proficient only with the mechanical manipulations is not sufficient for are successful completion of this course.

You will also be expected to develop and practice your verbal and written communication skills. You will be required to work in groups during class and will have ample opportunities to practice these skills. It is hoped that you will also find a group to work with outside of class.

The preface of the book (particularly page xi) is particularly well-written. (Pages vii – xi provide additional detail and insight into the methods you will encounter in this course.) Read the preface!

Use of Technology This course will make regular use of a graphing calculator. The TI-83 is the preferred calculator. A TI-82 is similar enough that it should be sufficient for our needs. (If you choose to use a different calculator, I do not promise to be able to assist you.) One of the objectives of this course is for you to become proficient in the appropriate use of your calculator. This includes recognizing when the calculator should be used and when hand manipulations are more efficient.

Course Content In this course we will be discussing most of the material from the first six chapters of the text. This will bring you into contact with the standard topics of calculus. That is, we will be looking at the way quantities change (differentiation) and accumulate (integration). We will also see several ways in which these concepts arise in other academic disciplines and everyday life.

Study Hints Before each class, you should both review the material from recent sections and read the section to be discussed that day. This will allow you to both understand my presentation of new material and identify questions that you need to resolve within your in-class group time.